# latexText page

### v1

@@ Abstract
Model-based iterative reconstruction plays a key role in solving inverse problems. However, the associated minimization problems are generally large-scale, ill-posed, nonsmooth, and sometimes even nonconvex, which present challenges in designing efficient iterative solvers and often prevent their practical use. Preconditioning methods can significantly accelerate the convergence of iterative methods. In some applications, computing preconditioners on-the-fly is beneficial. Moreover, forward models in image reconstruction are typically represented as operators, and the corresponding explicit matrices are often unavailable, which brings additional challenges in designing preconditioners. Therefore, for practical use, computing and applying preconditioners should be computationally inexpensive. This paper adapts the randomized Nyström approximation to compute effective preconditioners that accelerate image reconstruction without requiring an explicit matrix for the forward model. We leverage modern GPU computational platforms to compute the preconditioner on-the-fly. Moreover, we propose efficient approaches for applying the preconditioner to problems with nonsmooth regularizers. Our numerical results on image deblurring, super-resolution with impulsive noise, and computed tomography reconstruction demonstrate the efficiency and effectiveness of the proposed preconditioner.
@@ I Introduction
where $f(\bm{\mathrm{x}})$ denotes a data fidelity specifying the discrepancy between the model’s predictions and the degraded measurements. $g(\bm{\mathrm{x}})$ denotes a regularizer imposing prior information on $\bm{\mathrm{x}}$. $\lambda>0$ is a trade-off parameter to balance $f(\bm{\mathrm{x}})$ and $g(\bm{\mathrm{x}})$. Both $f(\bm{\mathrm{x}})$ and $g(\bm{\mathrm{x}})$ can be nonsmooth and nonconvex. This paper focuses on linear inverse problems where we solve
@@
where $\bm{\mathrm{A}}\in\mathbb{R}^{M\times N}$ with $M\leq N$ is called the forward model describing the mapping from the image of interest $\bm{\mathrm{x}}$ to the measurements $\bm{\mathrm{y}}$. In this paper, we assume $p\in(0,2]$.
In practice, the matrix $\bm{\mathrm{A}}$ is under-determined or ill-conditioned and the measurements are noisy, so one uses a regularizer $g(\bm{\mathrm{x}})$ to stabilize the solution. Typical choices of $g(\bm{\mathrm{x}})$ that are effective include wavelet [1], total variation (TV) [2], dictionary learning [3], low-rank [4]. When using such hand-crafted regularizers, we refer to (1) as variational image reconstruction. Over the past decades, deep learning (DL) [5] has received significant attention in the field of image reconstruction. Instead of using hand-crafted priors, DL-based methods learn implicit priors from massive training data. This includes end-to-end learning [6] and physics-informed deep unrolling [7, 8]. Alternatively, one can use plug-and-play [9, 10, 11] or regularization by denoising [12, 13] where one trains a deep denoiser [14] first and then uses it as an implicit prior for general inverse problems. Recently, building priors with generative models, such as diffusion models, has also attracted much attention. We refer the reader to [15] for a survey of those directions.
@@ I-A $\ell_{p}-\ell_{q}$ Image Reconstruction
which is known as the $\ell_{p}-\ell_{q}$ problem [16]. For $p,q\leq 1$, both $f(\bm{\mathrm{x}})$ and $g(\bm{\mathrm{x}})$ are nonsmooth. To address the nonsmooth challenge, Clason et al. considered (3) in its dual formulation [25], where the objective function becomes differentiable, at the cost of additional box and linear equality constraints. For general $p,\,q\in(0,2)$, Huang et al. [26] first smoothed $f$ and $g$ and then applied the majorization-minimization for the smoothed problem, leading to an inner least-squares problem. Then [26] used the Krylov based methods to solve the smoothed problem for acceleration. Following the same strategy, Buccini et al. [27] applied modulus-based iterative methods [28] to the smoothed problem to address box constraints. Alternatively, Chan et al. [29] solved (3) with a half-quadratic algorithm. Lanza et al. [30] proposed a Krylov subspace method to reduce computation of the method in [29] by assuming that the signal can be represented in a subspace.
Such solvers for (3) can be considered as iteratively reweighted methods (IRM) [16] because we eventually need to solve a linear equation (19) with updating of the associated weightings at each iteration. For completeness, Section III-A describes one type IRM for (3). Typically the appealing conjugate gradient (CG) method is chosen to address the associated linear equation. However, the convergence of CG can be extremely slow for some problems, especially for $p,\,q<1$, as the problem becomes nonconvex. Preconditioning methods [31, 32] are widely used to accelerate the convergence of CG. Because the weightings $\bm{\mathrm{W}}_{f}^{k}$ and $\bm{\mathrm{W}}_{g}^{f}$ (defined in (17)) change at each iteration, an effective preconditioner should change at each iteration, so computing such a preconditioner must be computationally inexpensive.
@@ I-B $\ell_{2}$ with Mixed Norm Image Reconstruction
When $p=2$, we consider $\Psi(\cdot)$ to be an $\ell_{1,\phi}$ mixed norm (i.e., $\|\mathcal{V}\|_{1,\phi}=\sum_{l=1}^{G}\|\bm{\mathrm{v}}_{l}\|_{\phi}$ with $\phi\geq 1$ and $\bm{\mathrm{v}}_{l}$ denoting the $l$th group of $\mathcal{V}=\{\bm{\mathrm{v}}_{1},\bm{\mathrm{v}}_{2},\cdots,\bm{\mathrm{v}}_{G}\}$), for which we can develop more efficient approaches than IRM to solve (3) and a broader range of choices for $\bm{\mathrm{L}}$. Moreover, we can also address convex constraints $\mathcal{C}$ easily. Dividing the components of $\mathcal{V}$ into $G$ groups, the mixed norm $\ell_{\psi,\phi}$ is defined by applying an $\ell_{\phi}$ norm to each group and then applying an $\ell_{\psi}$ norm to the entire set of groups. In this paper, we assume $\psi=1$. Specifically, we address
@@
Because (4) is a standard composite minimization problem, an appealing algorithm for solving it is the accelerated proximal gradient (APG) method [34]. Alternatively, one could use a primal-dual method [35, 36] or ADMM [37] etc. To accelerate the convergence of solving (4), one can adapt a quasi-Newton proximal method that uses second-order information [38, 39, 40]. However, the efficiency of using a quasi-Newton proximal method depends highly on the accuracy of approximating the Hessian matrix of $f(\bm{\mathrm{x}})$ and the efficiency of solving the associated weighted proximal mapping (WPM, defined in (5)).
@@ I-B1 Previous Work on Preconditioning Methods in Image Reconstruction
Preconditioning methods are used widely to accelerate the convergence of iterative methods [41] in scientific computing and have been extended to accelerate the convergence of image reconstruction. Clinthorne et al. [42] applied a spatially invariant preconditioner to accelerate the convergence of reconstructing single-photon emission computed tomography (SPECT) images with quadratic $g(\bm{\mathrm{x}})$. Fessler et al. [43] proposed a shift-variant preconditioner for accelerating positron emission tomography (PET) reconstruction with non-quadratic smooth $g(\bm{\mathrm{x}})$. Lin et al. [44] derived cosine transform and incomplete factorization based preconditioners for image super-resolution task with smooth $g(\bm{\mathrm{x}})$. Pelletier et al. [45] developed a block circulant with circulant block (BCCB) preconditioner for edge-preserving image super-resolution. The authors in [46, 47] considered diagonal preconditioners for accelerating PET and magnetic resonance imaging reconstruction. To reduce the computation of computing WPM, [48] discussed an unmatched preconditioner method for image reconstruction by computing WPM approximately. To address the nonsmoothness of $g(\bm{\mathrm{x}})$, Koolstra et al. [49] adapted the preconditioner to ADMM to accelerate the convergence of solving the associated least-squares problem. By leveraging deep learning, [50, 51] proposed methods to learn preconditioners to accelerate the convergence of image reconstruction.
@@ I-C Our Contribution
The preconditioners discussed in Section I-B1 were either used for smooth $g(\bm{\mathrm{x}})$ [42, 43, 44, 46, 47] or applied to a nonsmooth $g(\bm{\mathrm{x}})$ after smoothing [45, 51]. Moreover, the previous methods require domain knowledge because they were developed for specific applications, making them difficult to generalize to other applications. Although the preconditioners proposed in [49, 50] can be used to address nonsmooth $g(\bm{\mathrm{x}})$, the convergence rate of ADMM is only $\mathcal{O}(1/k)$. To address these challenges, we adapt the randomized Nyström approximation (RNA) to design an effective preconditioner, called randomized Nyström preconditioner (RNP), to accelerate variational image reconstruction. Our main contributions are summarized as follows.
[…]
  • By leveraging modern high-performance GPU computational platforms, we compute RNP on-the-fly, allowing it to accelerate IRM for (3) in a computationally efficient way. Our numerical experiments on image deblurring and super-resolution with impulsive noise showed that, by using RNP, we can reduce wall time by more than $90$%. To the best of our knowledge, this is the first work to consider the use of hardware platform in conjunction with computing preconditioners on-the-fly in image reconstruction.
@@ I-D Preliminaries and Roadmap
The rest of this paper is organized as follows. Section II describes how RNA can be a preconditioner for linear equations. Section III shows how to adapt RNP to accelerate IRM and APG for solving (3) and (4), respectively. Section IV presents numerical experiments on image deblurring and super-resolution with impulsive noise and computed tomography reconstruction to demonstrate the effectiveness and efficiency of using RNP for acceleration. Section V presets the conclusion and future work.
@@ II Randomized Nyström Approximation
The Nyström method provides a way to build a low-rank approximation for a symmetric positive semidefinite (SPSD) matrix $\bm{\Phi}\succeq 0\in\mathbb{R}^{N\times N}$. Let $\bm{\Omega}\in\mathbb{R}^{N\times K}$ be a random test matrix [54, 55] with sketch size $K\geq 1$. The Nyström method uses the following formula as a low-rank approximation of $\bm{\Phi}$, called the Nyström approximation
@@
where denotes the pseudo-inverse. Clearly $\bm{\Phi}\left<\bm{\Omega}\right>$ is also a SPSD matrix and its rank is at most $K$ [56, Lemma A.1]. Because a direct implementation of (9) is numerically unstable, we use the robust implementation proposed in [54], summarized in Algorithm 1. That algorithm computes $\bm{\mathrm{U}}$ and $\hat{\bm{\mathrm{S}}}$ such that $\bm{\Phi}\left<\bm{\Omega}\right>=\bm{\mathrm{U}}\hat{\bm{\mathrm{S}}}\bm{\mathrm{U}}^{\mathrm{T}}$, although in practice we do not compute that product explicitly.
@@
Theorem 2 shows that, by choosing an appropriate sketch size $K$, RNP can significantly reduce the condition number of $(\bm{\Phi}+\mu\bm{\mathrm{I}})$. Thus, one can significantly accelerate the convergence of iterative methods for solving (10) with RNP. However, in practice, computing $d_{\mathrm{eff}}$ is computationally expensive. Moreover, the value of $K=2\lceil 1.5d_{\mathrm{eff}}(\mu)+1\rceil$ can be very large, which will significantly increases the computation of computing and applying RNP. In our experimental settings, we found that using a moderate value for $K$ is sufficient. We refer the reader to [55, 54], where the authors discussed more theoretical properties of using RNP for considering (10). The main purpose of this paper is to study how to adapt RNP to accelerate variational image reconstruction, focusing on evaluating its effectiveness and efficiency. We also refer the reader to [55, 56, 58] and the references therein that discuss using RNP for other applications.
@@
0:  $\bm{\mathrm{U}}\in\mathbb{R}^{N\times K},~{}\hat{\bm{\mathrm{S}}}\in\mathbb{R}^{K\times K}$.
@@
2:  $\bm{\mathrm{Y}}=\bm{\Phi}\bm{\Omega}\in\mathbb{R}^{N\times K}$
3:  $v=\epsilon\|\bm{\Omega}\|_{F}$
4:  $\bm{\mathrm{Y}}_{v}=\bm{\mathrm{Y}}+v\bm{\Omega}$
5:  $\bm{\mathrm{C}}=\mathrm{chol}({\mathbf{\Omega}}^{\mathrm{T}}\bm{\mathrm{Y}}_{v})\in\mathbb{R}^{K\times K}$ Cholesky decomp.
6:  $\bm{\mathrm{B}}=\bm{\mathrm{Y}}_{v}\bm{\mathrm{C}}^{-{\mathrm{T}}}\in\mathbb{R}^{N\times K}$
7:  $[\bm{\mathrm{U}},\bm{\mathrm{S}},\sim]=\mathrm{SVD}(\bm{\mathrm{B}})$
8:  $\hat{\bm{\mathrm{S}}}=\max\{\bm{0},\bm{\mathrm{S}}^{2}-v\bm{\mathrm{I}}\}$

### v2

# Using Randomized Nyström Preconditioners to Accelerate Variational Image Reconstruction

@@ Abstract
Model-based iterative reconstruction plays a key role in solving inverse problems. However, the associated minimization problems are generally large-scale, nonsmooth, and sometimes even nonconvex, which present challenges in designing efficient iterative solvers. Preconditioning methods can significantly accelerate the convergence of iterative methods. In some applications, computing preconditioners on-the-fly is beneficial. Moreover, forward models in image reconstruction are typically represented as operators, and the corresponding explicit matrices are often unavailable, which brings additional challenges in designing preconditioners. Therefore, for practical use, computing and applying preconditioners should be computationally inexpensive. This paper adapts the randomized Nyström approximation to compute effective preconditioners that accelerate image reconstruction without requiring an explicit matrix for the forward model. We leverage modern GPU computational platforms to compute the preconditioner on-the-fly. Moreover, we propose efficient approaches for applying the preconditioners to problems with classical nonsmooth regularizers, i.e., wavelet, total variation, and Hessian Schatten-norm. Our numerical results on image deblurring, super-resolution with impulsive noise, and 2D computed tomography reconstruction illustrate the efficiency and effectiveness of the proposed preconditioner.
@@ I Introduction
where $f(\bm{\mathrm{x}})$ denotes a data-fidelity specifying the discrepancy between the model’s predictions and the degraded measurements, $g(\bm{\mathrm{x}})$ denotes a regularizer imposing prior information on $\bm{\mathrm{x}}$, and $\lambda>0$ is a trade-off parameter to balance $f(\bm{\mathrm{x}})$ and $g(\bm{\mathrm{x}})$. Both $f(\bm{\mathrm{x}})$ and $g(\bm{\mathrm{x}})$ can be nonsmooth and nonconvex. This paper focuses on linear inverse problems of the form
@@
where $\bm{\mathrm{A}}\in\mathbb{R}^{M\times N}$ with $M\leq N$ denotes the forward operator, which maps the unknown image $\bm{\mathrm{x}}$ to the measurements $\bm{\mathrm{y}}$. In this paper, we assume $p\in(0,2]$.
In practice, the matrix $\bm{\mathrm{A}}$ is under-determined or ill-conditioned and the measurements are noisy, so one uses a regularizer $g(\bm{\mathrm{x}})$ to stabilize the solution. Typical regularizer choices that are effective include wavelet sparsity [1], total variation (TV) [2], dictionary learning [3], and low-rank models [4]. When using such hand-crafted regularizers, we refer to (1) as variational image reconstruction. Over the past decades, deep learning (DL) [5] has received significant attention in the field of image reconstruction. Instead of using hand-crafted priors, DL-based methods learn implicit priors from massive training data. This includes end-to-end learning [6] and physics-informed deep unrolling [7, 8]. Alternatively, one can use plug-and-play (PnP) [9, 10, 11] or regularization by denoising (RED) [12, 13] where one trains a deep denoiser [14] first and then uses it as an implicit prior for general inverse problems. Recently, building priors with generative models, such as diffusion models, has also attracted much attention. We refer the reader to [15] for a survey of those directions.
@@ I-A $\ell_{p}-\ell_{q}$ Image Reconstruction
which is known as the $\ell_{p}-\ell_{q}$ problem [16]. For $p,q\leq 1$, both $f(\bm{\mathrm{x}})$ and $g(\bm{\mathrm{x}})$ are nonsmooth. To address the nonsmooth challenge, Clason et al. considered (3) in its dual formulation [25], where the objective function becomes differentiable, at the cost of additional box and linear equality constraints. For general $p,\,q\in(0,2)$, Huang et al. [26] first smoothed $f$ and $g$ and then applied the majorization-minimization for the smoothed problem, leading to an inner least-squares problem. Then, [26] used the Krylov subspace to represent the image, thereby reducing the dimension of the associated least-squares problem to that of the subspace. Consequently, it can be solved efficiently using a direct method. Following the same strategy, Buccini et al. [27] applied modulus-based iterative methods [28] to the smoothed problem to address the nonegativity constraint. Alternatively, Chan et al. [29] solved (3) with a half-quadratic algorithm. Lanza et al. [30] proposed a generalized Krylov subspace (GKS) method to reduce computation of the method in [29] by assuming that the signal can be represented in a subspace. Gazzola et al. [31] presented two flexible Krylov subspace methods for sparse reconstruction.
The solvers discussed above for (3) can be considered as iteratively reweighted methods (IRMs) [16] because we eventually need to solve a linear equation (19) with updating of the associated weightings at each iteration. For completeness, Section III-A describes one type IRM for (3). Typically the appealing conjugate gradient (CG) method is chosen to address the associated linear equation. However, the convergence of CG can be extremely slow for some problems, especially for p,q<1p,\,q<1, as the problem becomes nonconvex. Preconditioning methods [32, 33] are widely used to accelerate the convergence of CG. Because the weightings $\bm{\mathrm{W}}_{f}^{k}$ and $\bm{\mathrm{W}}_{g}^{f}$ (defined in (17)) change at each iteration, an effective preconditioner should change at each iteration, so computing such a preconditioner must be computationally inexpensive.
@@ I-B $\ell_{2}$ with Mixed Norm Image Reconstruction
When $p=2$, we consider $\Psi(\cdot)$ to be an $\ell_{1,\phi}$ mixed norm (i.e., $\|\mathcal{V}\|_{1,\phi}=\sum_{l=1}^{G}\|\bm{\mathrm{v}}_{l}\|_{\phi}$ with $\phi\geq 1$ and $\bm{\mathrm{v}}_{l}$ denoting the $l$th group of $\mathcal{V}=\{\bm{\mathrm{v}}_{1},\bm{\mathrm{v}}_{2},\cdots,\bm{\mathrm{v}}_{G}\}$), for which we can develop more efficient approaches than IRM to solve (3) and a broader range of choices for $\bm{\mathrm{L}}$. Moreover, we can also address convex constraints $\mathcal{C}$ easily. Dividing the components of $\mathcal{V}$ into $G$ groups, the mixed norm $\ell_{1,\phi}$ is defined by applying an $\ell_{\phi}$ norm to each group and then applying an $\ell_{1}$ norm to the entire set of groups. Specifically, we address
@@
Since (4) is a standard composite minimization problem, an appealing algorithm for solving it is the accelerated proximal gradient (APG) method [35]. Alternatively, one could use a primal-dual method [36, 37] or ADMM [38] etc. To accelerate the convergence of solving (4), one can adoapt a quasi-Newton proximal method that uses second-order information [39, 40, 41]. However, the efficiency of using a quasi-Newton proximal method depends highly on the accuracy of approximating the Hessian matrix of $f(\bm{\mathrm{x}})$ and the efficiency of solving the associated weighted proximal mapping (WPM, defined in (5)). For $\phi=2$, $g(\bm{\mathrm{x}})$ reduces to the group sparsity case, for which one can adapt a flexible Krylov subspace method [42] to solve (4).
@@ I-B1 Previous Work on Preconditioning Methods in Image Reconstruction
Preconditioning methods are used widely to accelerate the convergence of iterative methods [43] in scientific computing and have been extended to accelerate the convergence of image reconstruction. Clinthorne et al. [44] applied a spatially invariant preconditioner to accelerate the convergence of reconstructing single-photon emission computed tomography images with quadratic $g(\bm{\mathrm{x}})$. Fessler et al. [45] proposed a shift-variant preconditioner for accelerating positron emission tomography (PET) reconstruction with non-quadratic smooth $g(\bm{\mathrm{x}})$. Lin et al. [46] derived cosine transform and incomplete factorization based preconditioners for image super-resolution task with smooth $g(\bm{\mathrm{x}})$. Pelletier et al. [47] developed a block circulant with circulant block preconditioner for edge-preserving image super-resolution. The authors in [48, 49] considered diagonal preconditioners for accelerating PET and magnetic resonance imaging reconstruction. To reduce the computation of computing WPM, [50] discussed an unmatched preconditioner method for image reconstruction by computing WPM approximately. To address the nonsmoothness of $g(\bm{\mathrm{x}})$, Koolstra et al. [51] adapted the preconditioner to ADMM to accelerate the convergence of solving the associated least-squares problem. By leveraging deep learning, [52] proposed an approach to learn preconditioners to accelerate the convergence of image reconstruction.
@@ I-C Our Contribution
The preconditioners discussed in Section I-B1 were either used for smooth $g(\bm{\mathrm{x}})$ [44, 45, 46, 48, 49] or applied to a nonsmooth $g(\bm{\mathrm{x}})$ after smoothing [47]. Moreover, the previous methods require domain knowledge because they were developed for specific applications, making them difficult to generalize to other applications. Although the preconditioners proposed in [51, 52] can be used to address nonsmooth $g(\bm{\mathrm{x}})$, the convergence rate of ADMM is only $\mathcal{O}(1/k)$. To address these challenges, we adapt the randomized Nyström approximation (RNA) [53, 54] to design an effective preconditioner, called randomized Nyström preconditioner (RNP), to accelerate variational image reconstruction. Our main contributions are summarized as follows.
  • To the best of our knowledge, this is the first work to adapt RNP to accelerate image reconstruction with nonsmooth regularizers. Computing RNP requires only the evaluation of $\bm{\mathrm{A}}\bm{\mathrm{x}}$ and $\bm{\mathrm{A}}^{\mathrm{T}}\bm{\mathrm{x}}$ ($\mathrm{T}$} denotes the transpose operator), and does not require to access the explicit matrix $\bm{\mathrm{A}}$ or knowledge of any specific structure of $\bm{\mathrm{A}}$. Therefore, RNP can generalize to a wide range of applications.
  • By leveraging modern high-performance GPU computational platforms, we compute RNP on-the-fly, allowing it to accelerate IRM for (3) in a computationally efficient way. Our numerical experiments on image deblurring and super-resolution with impulsive noise showed that, by using RNP, we can reduce wall time by more than 9090%. To the best of our knowledge, this is the first work to consider the use of hardware platform in conjunction with computing preconditioners on-the-fly in image reconstruction.
@@ I-D Preliminaries and Roadmap
The rest of this paper is organized as follows. Section II describes how RNA can be a preconditioner for linear equations. Section III shows how to adapt RNP to accelerate IRM and APG for solving (3) and (4), respectively. Section IV presents numerical experiments on image deblurring and super-resolution with impulsive noise and computed tomography reconstruction to demonstrate the effectiveness and efficiency of using RNP for acceleration. Section V presents the conclusion and future work. Code to reproduce the results in the paper is available at https://github.com/hongtao-argmin/RNP-AccImageRecon.
@@ II Randomized Nyström Approximation
The Nyström method provides a way to build a low-rank approximation for a symmetric positive semidefinite (SPSD) matrix $\bm{\Phi}\succeq 0\in\mathbb{R}^{N\times N}$. Let $\bm{\Omega}\in\mathbb{R}^{N\times K}$ be an i.i.d. Gaussian random test matrix [53, 54] with sketch size $K\geq 1$. The Nyström method uses the following formula as a low-rank approximation of $\bm{\Phi}$, called the Nyström approximation
@@
where denotes the pseudo-inverse. Clearly $\bm{\Phi}\left<\bm{\Omega}\right$> is also a SPSD matrix and its rank is at most $K$ [57, Lemma A.1]. Because a direct implementation of (9) is numerically unstable, we adopt the robust scheme proposed in [53], which is summarized in Algorithm 1. This algorithm computes $\bm{\mathrm{U}}$ and $\hat{\bm{\mathrm{S}}}$ such that $\bm{\Phi}\langle\bm{\Omega}\rangle=\bm{\mathrm{U}}\hat{\bm{\mathrm{S}}}\bm{\mathrm{U}}^{\mathrm{T}}$, although this product is not formed explicitly in practice.
@@
Theorem 2 shows that, by choosing an appropriate sketch size $K$, RNP can significantly reduce the condition number of $(\bm{\Phi}+\mu\bm{\mathrm{I}})$. Thus, one can significantly accelerate the convergence of iterative methods for solving (10) with RNP. However, in practice, computing $d_{\mathrm{eff}}$ is computationally expensive. Moreover, the value of $K=2\lceil 1.5d_{\mathrm{eff}}(\mu)+1\rceil$ can be very large, which will significantly increase the computation of computing and applying RNP. In our experimental settings, we found that using a moderate value for $K$ is sufficient. We refer the reader to [54, 53], where the authors discussed more theoretical properties of using RNP for considering (10). The main purpose of this paper is to study how to adapt RNP to accelerate variational image reconstruction, focusing on evaluating its effectiveness and efficiency. We also refer the reader to [54, 57, 59] and the references therein that discuss using RNP for other applications. Although Theorem 2 presents the split preconditioner $\bm{\mathrm{P}}^{-1/2}$, we use $\bm{\mathrm{P}}^{-1}$ in practice for simplicity.
@@
[…]
@@
[…]